OpenStudy (anonymous):

A scuba diver descended 16 5/12 feet below sea level. Then, he descended another 3 3/5 feet. Which of the following is true about the scuba diver after both descents? A. The location of the scuba diver in relation to sea level was -12 46/60 feet. B. The location of the scuba diver in relation to sea level was 12 46/60 feet. C. The location of the scuba diver in relation to sea level was 20 1/6 feet. D. The location of the scuba diver in relation to sea level was -20 1/6 feet.

OpenStudy (anonymous):

Let me do some work and I will get back to you?

OpenStudy (anonymous):

ok @cherry18

OpenStudy (anonymous):

you are trying to find distance. Distance is always POSITIVE. So A and D is out.

OpenStudy (anonymous):

Can you select more than 1?

OpenStudy (anonymous):

no, just 1

OpenStudy (anonymous):

Ok. First you have to add the numbers. So what is the common# between 12 and 5?

OpenStudy (anonymous):

60

OpenStudy (anonymous):

Good.

OpenStudy (anonymous):

So that is the common denominater. Now finish the problem and tell me the final answer you get.

OpenStudy (anonymous):

@Freya_Irie

OpenStudy (anonymous):

Yeah im here, OS is lagging.

OpenStudy (anonymous):

You mean ols? k12

OpenStudy (anonymous):

no i mean OS openstudy, im lagging over here.

OpenStudy (anonymous):

oh ok.

OpenStudy (anonymous):

Any way, i would go with B

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

That is incorrect. Sorry. It is C.

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

Good try. If you finish the problem you get answer C.

OpenStudy (anonymous):

ok, can u help with 1 more?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

Oh. Thanks for the medal by the way.

OpenStudy (anonymous):

J & B Research is conducting a survey to see if people still live in the state in which they were born. They hired Kimberly, Emily, and Ava to conduct the survey. Kimberly surveyed 400 people with 38% of respondents answering yes. Emily surveyed 364 people with 3 out of 7 respondents answering yes. Ava surveyed 350 people with 0.44 of respondents answering yes. Place the survey takers in order from least number of respondents answering yes to greatest number of respondents answering yes. _______, _______, _______

OpenStudy (anonymous):

Ok. Ava, Kim, Emily

OpenStudy (anonymous):

Thanks...

OpenStudy (anonymous):

Welcome!....

OpenStudy (anonymous):

D. is very close. The answer is:\[-\left(3+\frac{3}{5}\right)-\left(16+\frac{5}{12}\right)=-\frac{1201}{60}=-\left(20+\frac{1}{60}\right) \]

OpenStudy (anonymous):

So. C. I was right.